The Fashion Shoe Company operates a chain of women’s shoe shops that carry many styles of shoes that are all sold at the same price. Sales personnel in the shops are paid a sales commission on each pair of shoes sold plus a small base salary. The following data pertains to Shop 48 and is typical of the company’s many outlets: Per Pair of Shoes Selling price $ 40.00 Variable expenses: Invoice cost $ 19.50 Sales commission 4.50 Total variable expenses $ 24.00 Annual Fixed expenses: Advertising $ 38,000 Rent 28,000 Salaries 140,000 Total fixed expenses $ 206,000 Problem 5-26 Part 5 5. Refer to the original data. As an alternative to (4) above, the company is considering paying the Shop 48 store manager 55 cents commission on each pair of shoes sold in excess of the break-even point. If this change is made, what will be Shop 48's net operating income (loss) if 16,175 pairs of shoes are sold?
Answer #1
Break even point in pairs of shoes = 206000/(40-24) = 12875 Pairs
5] Sales = 16175*40 = $ 6,47,000.00
Less: Variable cost = 16175*24 = $ 3,88,200.00
Less: Commission to stores manager = (16175-12875)*0.55 = $        1,815.00
Less: Fixed costs $    2,06,000.00
Net operating income $      50,985.00